• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name B. Quinn

    Log entry time 08:58:00 on October15,2009

    Entry number 295029

    keyword=Lumi widths

    Lumi widths were large (~1500ppm) at start of run 13280.  After about 15000 
    pairs beam current was lowered from ~99.5uA to ~96uA.  Widths decreased but are 
    still a little wide (1100-1200 instead of <1100).
    ...improved further as I was typing this... see screen shot.

    FIGURE 1